OpenStudy (anonymous):

Given the parent functions f(x) = log2 (3x − 9) and g(x) = log2 (x − 3), what is f(x) − g(x)?

OpenStudy (anonymous):

@e.mccormick

OpenStudy (mathstudent55):

Subtract the functions.

OpenStudy (mathstudent55):

\(\Large f(x) = \log_2 (3x - 9)\) \(\Large g(x) = \log_2 (x - 3)\) \(\Large f(x) - g(x) = \log_2 (3x - 9) - \log_2 (x - 3)\) Ok so far?

OpenStudy (anonymous):

yes @mathstudent55

OpenStudy (mathstudent55):

Now we need to simplify the fright side using some rules of logs. Here are two rules of logs: \(\Large \log ab = \log a + \log b\) \(\Large \log \dfrac{a}{b} = \log a - \log b\)

OpenStudy (anonymous):

Alright.

OpenStudy (anonymous):

Ok, i thought we were going to factor lol

OpenStudy (mathstudent55):

Yes, we do factor. Look at the first log below compared to what it was. \(\Large f(x) - g(x) = \log_2 3(x - 3) - \log_2 (x - 3)\)

OpenStudy (anonymous):

you moved the 3 out of (3x-9) so u divide by 3

OpenStudy (mathstudent55):

I only factored out the 3. 3(x - 3) = 3x - 9. We are still taking the log of the same quantity, only it is now written in a different form, its factored form. Now we use the first rule of logs above to deal with the first log in our equation.

OpenStudy (mathstudent55):

I'm now working on the red part: \(\Large f(x) - g(x) = \color{red}{\log_2 3(x - 3)} - \log_2 (x - 3)\) \(\Large f(x) - g(x) = \color{red}{\log_2 3 + \log_2 (x - 3)} - \log_2 (x - 3)\) You see how the rule of logs is applied?

OpenStudy (anonymous):

Yes :)

OpenStudy (mathstudent55):

What can you do to the green part below? \(\Large f(x) - g(x) = \log_2 3 + \color{green}{\log(x - 3) - \log_2 (x - 3)}\)

OpenStudy (anonymous):

Cancel out?

OpenStudy (anonymous):

would the answer be log 2 1/3

OpenStudy (mathstudent55):

Correct.

OpenStudy (mathstudent55):

\(\Large f(x) - g(x) = \log_2 3 \)

OpenStudy (anonymous):

is it log 2 3 or log 2 1/3?

OpenStudy (mathstudent55):

Answer is: \(\Large f(x) - g(x) = \log_2 3 \)

OpenStudy (anonymous):

Alright, Thanksss! :)

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (anonymous):

Given the parent functions f(x) = log10 x and g(x) = 5x − 2, what is f(x) • g(x)? for this one, i got f(x) • g(x) = log10 x5x − 2

OpenStudy (anonymous):

Was I right?

OpenStudy (mathstudent55):

Here you need to multiply a log by a binomial. \(\Large f(x) = \log_{10} x\) \(\Large g(x) = 5x − 2\) Multiply the left sides together and multiply the right sides together: \(\Large f(x) \cdot g(x) = \log_{10} x \cdot (5x - 2)\) Use the commutative property on the right side: \(\Large \color{red}{f(x) \cdot g(x) = (5x - 2)\log_{10} x} \) Now use the distributive property: \(\Large \color{green}{f(x) \cdot g(x) = 5x\log_{10} x - 2 \log_{10} x}\)

OpenStudy (mathstudent55):

You were close. The idea was correct, but you need to use parentheses.

OpenStudy (mathstudent55):

Any of the lines above is a correct answer, but to make the line in black clearer, I think it's better to write it as: \(\Large f(x) \cdot g(x) = (\log_{10} x) (5x - 2)\)

OpenStudy (anonymous):

Okay! Thanks :)